A line is perpendicular to 5x+2y-8=0 and has the same y intercept as x+4y-12=0. Find the equation for the line.
Answer by Boreal(15235)

same y-intercept means (0, y) are the same for both. The intercept of the second equation is found by writing it in slope intercept form, and it is 4y=-x+12 or y=-(1/4)x + 3, so the intercept is (0, 3)
2y=-5x+8 for the first equation or y=-(5/2)x+4 in slope-intercept form.
Perpendicular lines had a negative reciprocal for the slope product of both.
The line desired has slope (+2/5) with y-intercept 3 and its equation is y=(2/5)x+3 ANSWER